OpenStudy (anonymous):

verify the Identity cos4u=cos^(2)(2u) - sin^(2)(2u)

OpenStudy (cwrw238):

compare with standard identity cos(2x) = cos^2(x) - sin^2(x)

OpenStudy (cwrw238):

replace x with 2u

OpenStudy (anonymous):

@cwrw238 so what is the answer?

zepdrix (zepdrix):

Oh I thought those were 22's lolol, I'm glad you fixed it ambition :D I was really scratching my head over this one earlier.

zepdrix (zepdrix):

As penguin guy said, this is the Double Angle Formula for Cosine. \[\huge \cos (2u)=\cos^2u-\sin^2u\] Let's apply this to our problem,\[\large \cos(4u) \quad = \quad \cos(2\cdot 2u)\]\[\large \cos(2\cdot 2u) \quad = \quad \cos^2(2u)-\sin^2(2u)\]

OpenStudy (anonymous):

ok so is there something else i have to do?

OpenStudy (anonymous):

@zeepdrix

OpenStudy (anonymous):

@zepdrix

OpenStudy (cwrw238):

no - xepdrix as done it all for you